OpenStudy (zuha):

Which ratios form a proportion? (1 point) 4/9, 12 / 25 3 / 7, 18 / 42 5/11, 20/45 5/8, 15/25 2. Which ratio forms a proportion with 18 over 30? (1 point) 20 / 35 50 / 90 36 / 48 60 / 100 3. Which proportion has cross products of 5 × 24 and 8 × 15? (1 point) a 5/8=15/24 b 5/24=8/15 c 8/24=15/5 d 15/8=5/24

OpenStudy (anonymous):

For question one just see which ratio pair differ by a common multiple. We can see that B is the correct answer. For question two do the same thing however you want to equate each answer ratio to 18/30 and see if any differ by simply a factor of something. We can see that the answer is D. For question three cross products are simply the denominator of ratio A and the Numerator for ratio B multiplied and the same for the second pair of numbers left seeing that we can see that A is the correct answer to the question.

OpenStudy (zuha):

Thank you so much for answering with explanation , you didn't give me straight answer but explained it as well :)

OpenStudy (zuha):

Please can you explain more how you picked the answer for question 2

OpenStudy (anonymous):

You simply want to see what ratio of the provided answers is equivalent to 18/30. So the setup should be like this: \[\frac{ 18 }{30 }=\frac{ x }{ y }....18y=30x....\] Or you can just use the answers to see which one will simplify to the simpliest form of each which is answer D. Both Reduce down to \[\frac{ 3 }{ 5 }\]

OpenStudy (zuha):

thanks got it :)